Given that f(x) =x^2-3 and g(x)=2x+1 find
1)(fg)(0)
2)(f/g)(-squareroot3)

Answer by stanbon(75887)

Given that f(x) =x^2-3 and g(x)=2x+1 find
1)(fg)(0) = f[g(0)] = f[2*0+1) = f(1) = 1^2-3 = -2
======================

2)(f/g)(-squareroot3) = [f(-sqrt(3)]/g[(-sqrt(3)]
---
= (-sqrt(3))^2-3 / [2(-sqrt(3))+1]
----
= (3-3)/[-2sqrt(3)+1]
---
= 0
============
Cheers,
Stan H.